# Sublinear Data Structures for Nearest Neighbor in Ultra High Dimensions

### v1

@@ 5.2.4 Space and Query Complexity
$\left((\sigma_{1}(U_{i,1}-U_{j,1}))^{2}+\cdots+(\sigma_{n}(U_{i,n}-U_{j,n}))^{2}\right)\left(\sum_{k\in[n]}(V_{b,k})^{2}\right)\geq(\sigma_{1}V_{1,b}(U_{i,1}-U_{j,1})+\cdots+\sigma_{n}V_{n,b}(U_{i,n}-U_{j,n}))^{2}$
which proves (2). ∎
The proof of the following claim is analogous to the proof of 5.15 and is omitted for succinctness.
Claim 5.28.
@@
Lemma 5.29.
With probability $1-\delta^{\prime}$, the space complexity of Algorithm 4 is $O(\frac{n\log^{2}(n/\delta)\log(d)}{\epsilon^{6}\delta^{4}})$ wordsize, and the number of coordinates that we query is $O(\frac{n\log(n/\delta)}{\epsilon^{4}\delta^{2}})$.
Proof.
By 5.28, we have that the number of coordinates of $q$ that we query is $|I|=O(Tn)=O(\frac{n\log(n/\delta)}{\epsilon^{4}\delta^{2}})$. Next we analyze the space complexity of the data structure. Matrix $M$ is of wordsize $O(m|I|)$, $M(R)$ is of wordsize $O(mn)$, $I$ is of wordsize $O(|I|\log(d))$. Thus the space complexity of the our data structure is $O(m|I|+mn+|I|\log(d))=O(\frac{n\log^{2}(n/\delta)\log(d)}{\epsilon^{6}\delta^{4}})$ wordsize. ∎
@@
6 Data Structure with $\tilde{O}(n^{2})$ Space for General $\ell_{p}$ Metrics
@@
In this section, we describe our data structure with space $O(n^{2}\log(n/\delta)\log(d)(\log\log n)^{p}/\epsilon^{p+2})$ and query time $O(n\log(n/\delta)(\log\log n)^{p}/\epsilon^{p+2})$ for the approximate nearest neighbor problem for the $\ell_{p}$ metric for any $p$ and prove Theorem 6.1.
Theorem 6.1.
@@
6.1 A Data Structure for Two Input Points
We start with a data structure (see Algorithm 5) answering approximate nearest neighbor queries for the case of $2$ points, say $a$ and $b$. It generates Theorem 4.2, and the main differences are: (1) It works for $\ell_{p}$ metric. (2) It explicitly use truncation to remove the bounding box assumption.
Lemma 6.2.
@@ 17            Truncate $q^{\prime(i)}\leftarrow\mathsf{truncate}_{\{l^{(i)},u^{(i)}\}}(q^{(i)})$;
Claim 6.4.
(Truncation preserves ratio) For $\ell_{p}$ metric, let $q^{\prime}$ be a truncated point, where $q^{\prime(i)}=\mathsf{truncate}_{\{l^{(i)},u^{(i)}\}}(q^{(i)})$. $l^{(i)},u^{(i)}$ are defined in Line 5 and 5 of Algorithm 5.
@@
Proof.
(For Claim 6.4) Recall our assumption that $a=\mathbbm{0}$ and $\forall i,b_{i}\geq 0$. We prove that if $\mathsf{dist}(a,q)/\mathsf{dist}(b,q)\geq 1+\epsilon$,
@@
Case 2: $q^{(1)}\leq q^{\prime(1)}\leq 0$, the argument is similar. ∎
6.2 A Data Structure for $n$ Points
In this section, we prove Theorem 6.1. Let $C=\{c_{1},c_{2},\dots c_{n}\}\subseteq\mathbb{R}^{d}$ denote the set of input points and let $q\in\mathbb{R}^{d}$ denote the query point. Our goal is to find an approximate nearest neighbor of $q$ in the set $C$. The argument used here is from [AFHN16], but we still present it to be self-contained.
@@ 6.2 A Data Structure for $n$ Points
For all $(c_{i},c_{j})\in\binom{C}{2}$, we run the preprocessing steps of the algorithm guaranteed by Lemma 6.2, where we set $a\leftarrow c_{i}$ and $b\leftarrow c_{j}$.
@@
We first cast the nearest neighbor problem in the language of finding the minimum value in an array of length $n$ using only imperfect comparison queries as follows. Consider an array of length $n$ whose values are $\mathsf{dist}(q,c_{i})$ for $i\in[n]$. Although we do not have access to these values directly, Lemma 6.2 guarantees an algorithm that, given two indices $i,j\in[n]$, with probability at least $1-\delta$, returns the symbol $>$ if $\mathsf{dist}(q,c_{i})>(1+\epsilon)\mathsf{dist}(q,c_{j})$ and returns the symbol $<$ if $(1+\epsilon)\mathsf{dist}(q,c_{i})<\mathsf{dist}(q,c_{j})$. In other words, we have access to a comparison operator for the entries of the array such that the comparator works correctly with high probability, if the entries are more than a factor of $1+\epsilon$ apart. We use the following recursive algorithm, called initially with $S$ set to $[n]$, can be used to find an approximate minimum in this case.
@@
    If $|S|=2$, run the algorithm guaranteed by Lemma 6.2 and return the index of the smaller value. Note that this step corresponds to the use of the imperfect comparison operator.
@@
      For each pair $(m_{i},m_{j})\in\binom{M}{2}$, run the algorithm guaranteed by Lemma 6.2.
@@
The recursion depth of the algorithm is $\log\log n$ by virtue of the recurrence $t(n)=t(\sqrt{n})+1$. The total number of calls to the query algorithm of Lemma 6.2 at each recursion level is $\Theta(n)$ and hence, overall, this 2-point query algorithm is invoked $O(n\log\log n)$ times.
@@
We first condition on the good event that every call to the query algorithm guaranteed by Lemma 6.2 is correct. This event happens with probability at least $1-\Theta(n\log\log n)\cdot\delta$. Next, we argue about the deterioration in the approximation guarantee. The main idea is that Step 2(d) and 2(e) together return a $(1+\epsilon)^{2}$-approximate nearest neighbor of $q$ in the set $M$, given our conditioning. Let $m^{*}\in M$ denote the nearest neighbor of $q$ in $M$. Assume for the sake of contradiction that the point $m\in M$ returned in Step 2(e) is such that $\mathsf{dist}(q,m)>(1+\epsilon)^{2}\cdot\mathsf{dist}(q,m^{*})$. By the guarantee of our imperfect comparison operator, $m$ does not ”win” against $m^{*}$ in the 2-point comparison game. Consider now an $m^{\prime\prime}\in M$ that $m$ wins against in the 2-point comparison game. If $\mathsf{dist}(m^{\prime\prime},q)>(1+\epsilon)\cdot\mathsf{dist}(m,q)$, then it must also be the case that $m^{*}$ wins against $m^{\prime\prime}$ in the comparison game. Even if $\mathsf{dist}(m^{\prime\prime},q)\leq(1+\epsilon)\cdot\mathsf{dist}(m,q)$ and $m$ wins against $m^{\prime\prime}$ because of the do not care condition, the distance of $m^{\prime\prime}$ to $q$ is still a $(1+\epsilon)$ factor more than the distance of $m^{*}$ to $q$. Thus, $m^{*}$ one more 2-point comparison game as compared to $m$ and this is a contradiction. Since we incur a loss of a factor of $(1+\epsilon)^{2}$ in each recursion level, overall, the loss in approximation guarantee is $(1+\epsilon)^{2\log\log n}$. ∎
Proof of Theorem 6.1.
@@
To get our data structure, we run each invocation of the algorithm of Lemma 6.2 above (i.e., in Min-Finding) with $\epsilon=\epsilon^{\prime}/\log\log n$ and $\delta=\delta^{\prime}/\Theta(n\log\log n)$, we can get a $(1+\epsilon^{\prime})$-approximate nearest neighbor data structure with success probability at least $1-\delta^{\prime}$.

### v2

@@ B.2.4 Space and Query Complexity
$\left((\sigma_{1}(U_{i,1}-U_{j,1}))^{2}+\cdots+(\sigma_{n}(U_{i,n}-U_{j,n}))^{2}\right)\left(\sum_{k\in[n]}(V_{b,k})^{2}\right)$
$\geq(\sigma_{1}V_{1,b}(U_{i,1}-U_{j,1})+\cdots+\sigma_{n}V_{n,b}(U_{i,n}-U_{j,n}))^{2}$
which proves (2).
The proof of the following claim is analogous to the proof of 8 and is omitted for succinctness.
Claim 14.
@@
Lemma B.21.
With prob. $1-\delta^{\prime}$, the space complexity of Algorithm 4 is $O(\frac{n\log^{2}(n/\delta)\log(d)}{\epsilon^{6}\delta^{4}})$ wordsize, and the number of coordinates that we query is $O(\frac{n\log(n/\delta)}{\epsilon^{4}\delta^{2}})$.
Proof B.22.
By 14, we have that the number of coordinates of $q$ that we query is $|I|=O(Tn)=O(\frac{n\log(n/\delta)}{\epsilon^{4}\delta^{2}})$. Next we analyze the space complexity of the data structure. Matrix $M$ is of wordsize $O(m|I|)$, $M(R)$ is of wordsize $O(mn)$, $I$ is of wordsize $O(|I|\log(d))$. Thus the space complexity of the our data structure is $O(m|I|+mn+|I|\log(d))=O(\frac{n\log^{2}(n/\delta)\log(d)}{\epsilon^{6}\delta^{4}})$ wordsize.
@@
Appendix C $O(1)$-Approximate Sublinear Representation of Clusterings
Theorem C.1.
Definition C.2 (Cost of clustering).
Consider a set $C=\{c_{1},\dots,c_{k}\}$ of $k$ centers, and a set $P$ of points in $\mathbb{R}^{d}$ equipped with $\ell_{p}$-metric. The cost of $C$ w.r.t. partition $\sigma:P\rightarrow[k]$ is $\mathsf{cost}_{p}(C,P,\sigma)=\sum_{q\in P}\|q-\sigma(q)\|_{p}^{p}$. The cost $\mathsf{cost}_{p}(C,P)=\sum_{q\in P}\min_{c_{i}\in C}\|q-c_{i}\|_{p}^{p}$ of $C$ is defined as the minimum cost of any partitioning. We call $\mathsf{cost}_{1}(C,P)$ the $k$-Median cost and $\mathsf{cost}_{2}(C,P)$ the $k$-Means cost of $C$, respectively.
Corollary C.3.
Consider a set $C$ of $k$ centers $c_{1},\dots,c_{k}\in[-\Delta,\Delta]^{d}$ equipped with $\ell_{p}$ metric ($p=1,2$). We can efficiently construct a randomized data structure such that for any set $P\subset[-\Delta,\Delta]^{d}$ of points the following conditions hold. (i) The data structure reads $O(k\mathrm{polylog}(k,d,\Delta)/(\epsilon^{3}))$ coordinates of $P$. (ii) It returns a partition $\sigma:P\rightarrow[k]$ (iii) $\mathsf{cost}_{p}(C,P,\sigma)\leq O(1)\mathsf{cost}_{p}(C,P)$. (iv) The data structure has a size of $O(k\mathrm{polylog}(k,d,\Delta)/(\epsilon^{3}))$ words.
3
4      for $t\in[T]$ do
5             for $b\in[d]$ do
6                   Add $b$ to $I$ with probability $p^{(b)}\triangleq\max_{(i,j)\in{\binom{n}{2}}}\frac{\left|c_{i}^{(b)}-c_{j}^{(b)}\right|}{\|c_{i}-c_{j}\|_{1}}$;
7
8
10       Let $m\leftarrow O(\log(n)\log(\Delta d)/(\epsilon^{2}))$;
11       Sample $M\in\mathbb{R}^{[m]\times I}$, where each entry of $M$ follows the Cauchy distribution, whose density function is $c(x)=\frac{1}{\pi(1+x^{2})}$. Notice that $M(\cdot)\colon\mathbb{R}^{I}\rightarrow\mathbb{R}^{m}$ is an oblivious linear mapping [23];
12
13      store $I$, $M(R)=\{Mr_{i}|i\in[n]\}$, $\{p^{(b)}|b\in I\}$, $M$
14Query $(q)$ // Inputs: $q\in[-\Delta,\Delta]^{d}$
15       Query $q^{(b)},\forall b\in I$;
16       Let $u\in\mathbb{R}^{I}$, where $\forall b\in I,u^{(b)}=q^{(b)}/p^{(b)}$;
18       Let $\hat{i}=\operatorname*{arg\,min}_{i\in[n]}F(Mr_{i},Mu)$;
19       return $\hat{i}$
Theorem C.4.
$\mathbb{E}[\Gamma]=O(1)$.
Lemma C.5.
It holds that $\mathbb{P}[G_{1}]\geq 1-\epsilon/(4\Delta d)$
Definition C.6 (Good and bad regions).
Consider two points $c_{1},c_{2}\in\mathbb{R}^{d}$, and a query point $q\in\mathbb{R}^{d}$. We define the good region $\mathsf{Good}(c_{1},c_{2},q)$ w.r.t. $(c_{1},c_{2},q)$ as
$\{z\in[d]\mid\min\{c_{1}^{(z)},c_{2}^{(z)}\}-\frac{8}{\epsilon}|c_{1}^{(z)}-c_{2}^{(z)}|\leq q^{(z)}\leq\max\{c_{1}^{(z)},c_{2}^{(z)}\}+\frac{8}{\epsilon}|c_{1}^{(z)}-c_{2}^{(z)}|\}$
Lemma C.7.
$\mathbb{P}[G_{2}]\geq 1-\epsilon/(4\Delta d)$
Proof C.8.
Using Chernoff bound,
$\Pr\left[\left|\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}-\mathbb{E}\left[\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}\right]\right|\geq\epsilon_{2}\mathbb{E}\left[\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}\right]\right]$
$\leq$ $2\exp\left(-\frac{\epsilon_{2}^{2}T\mathbb{E}[\sum_{z\in[d]}P_{t,z}]}{3(1+\frac{8}{\epsilon})\|c_{i}-c_{i^{*}}\|_{1}}\right)$
$=$ $2\exp\left(-\frac{\epsilon_{2}^{2}T}{3(1+\frac{8}{\epsilon})}\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},q)}{\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{i^{*}})}\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{i^{*}})}{\|c_{i}-c_{i^{*}}\|_{1}}\right)$
$\leq$ $2\exp\left(-\frac{\epsilon_{2}^{2}T}{3(1+\frac{8}{\epsilon})}\cdot\frac{1}{2}\cdot\frac{13}{16}\right)\leq\frac{1}{2\Delta n},$
We conclude the lemma by union bound over all $c_{i}\neq c_{i^{*}}$ such that $(\|c_{i}-q\|_{1})/(\|c_{i^{*}}-q\|_{1})\geq 1+\epsilon$.
Let $G=G_{1}\wedge G_{2}$. Lemmas C.5 and C.7, together with union bound gives the following.
Lemma C.9.
$\mathbb{P}[G]\geq 1-\epsilon/(2\Delta d)$.
Lemma C.10.
Proof C.11.
Since we know that $\mathsf{dist}_{\mathsf{Bad}}(c_{i},q)-(1+\epsilon)\mathsf{dist}_{\mathsf{Bad}}(c_{i^{*}},q)\leq 0$, we have $\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},q)}{\mathsf{dist}_{\mathsf{Good}}(c_{i^{*}},q)+\mathsf{dist}_{\mathsf{Bad}}(c_{i^{*}},q)}\geq 2\max\{1,\lambda\}$, which is $\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},q)}{\|c_{i^{*}}-q\|_{1}}\geq 2\max\{1,\lambda\}$. Since $\mathsf{dist}_{\mathsf{Good}}(r_{i},u)\approx_{\epsilon}T\cdot\mathsf{dist}_{\mathsf{Good}}(c_{i},q)$, and $\|r_{i^{*}}-u\|_{1}=\lambda\cdot T\cdot\|c_{i^{*}}-q\|_{1}$, we have $\frac{\mathsf{dist}_{\mathsf{Good}}(r_{i},u)}{\|r_{i^{*}}-u\|_{1}}\geq\frac{2\max\{1,\lambda\}}{(1+\epsilon)\lambda}$. Thus $\frac{\|r_{i}-u\|_{1}}{\|r_{i^{*}}-u\|_{1}}\geq 2/(1+\epsilon)\geq(1+\epsilon)^{2}$.
We next bound the probability that the algorithm outputs a far center.
Lemma C.12.
Proof C.13.
By Lemma C.10 it follows that $\mathsf{dist}(r_{i},u)>(1+\epsilon)^{2}\mathsf{dist}(r_{i^{*}},u)$. Since the event $G$ contains the event $G_{1}$ it follows that $F(Mr_{i},Mu)>F(Mr_{i^{*}},Mu)$. Thus, the algorithm does not output $i$.
Lemma C.14.
Proof C.15.
$\leq 4\max\{\lambda,1\}$
The last inequality follows from Lemma C.12.
We use Lemma C.14 to conclude the proof for the constant approximation factor.
Proof C.16 (Proof of C.4).
$\leq\epsilon+4\mathbb{E}[\Lambda]+4\leq 8+\epsilon$ (6)
In Line 3 we use the fact that the maximal distance inside the $[-\Delta,\Delta]^{d}$ grid is $2\Delta d$ and Lemma C.9. In Line 4 we use the fact $\mathbb{P}[\Lambda=\lambda\mid G]\leq\mathbb{P}[\Lambda=\lambda]/\mathbb{P}[G]$. In Line 5 we use Lemma C.14. In Line 6 we use the fact that $\mathbb{E}[\Lambda]=1$.
Claim 15.
With probability $1-1/(nd\Delta)$, we have $|I|\leq 2Tn$.
Proof C.17.
Using Chernoff bound, we have
Thus, with probability $1/(nd\Delta)$, we have $|I|\leq 2E[|I|]\leq 2Tn$.
Lemma C.18.
Proof C.19.
By 15, we have that the number of coordinates of $q$ that we query is $|I|=O(Tn)=O(\frac{n\log(n)\log(\Delta d)}{\epsilon^{3}})$. Next we analyze the space complexity of the data structure. Matrix $M$ is of wordsize $O(m|I|)$, $M(R)$ is of wordsize $O(mn)$, $I$ is of wordsize $O(|I|\log(d))$. Thus the space complexity of the our data strcuture is $O(m|I|+mn+|I|\log(d))=O(n\mathrm{polylog}(n,d,\Delta)/(\epsilon^{3}))$ wordsize.
@@ 18       Let $\hat{i}=\operatorname*{arg\,min}_{i\in[n]}F(Mr_{i},Mu)$;
Appendix D Data Structure with $\tilde{O}(n^{2})$ Space for General $\ell_{p}$ Metrics
@@
In this section, we describe our data structure with $O(n^{2}\log(n/\delta)\log(d)(\log\log n)^{p}/\epsilon^{p+2})$ space and $O(n\log(n/\delta)(\log\log n)^{p}/\epsilon^{p+2})$ query time for the approximate nearest neighbor problem for the $\ell_{p}$ metric for any $p$ and prove Theorem D.1.
Theorem D.1.
@@
D.1 A Data Structure for Two Input Points
We start with a data structure (see Algorithm 6) answering approximate nearest neighbor queries for the case of $2$ points, say $a$ and $b$. It generates Theorem A.2, and the main differences are: (1) It works for $\ell_{p}$ metric. (2) It explicitly use truncation to remove the bounding box assumption.
Lemma D.2.
@@ 17            Truncate $q^{\prime(i)}\leftarrow\mathsf{truncate}_{\{l^{(i)},u^{(i)}\}}(q^{(i)})$;
Claim 16.
(Truncation preserves ratio) For $\ell_{p}$ metric, let $q^{\prime}$ be a truncated point, where $q^{\prime(i)}=\mathsf{truncate}_{\{l^{(i)},u^{(i)}\}}(q^{(i)})$. $l^{(i)},u^{(i)}$ are defined in Line 6 and 6 of Algorithm 6.
@@
Proof D.5.
(For Claim 16) Recall our assumption that $a=\mathbbm{0}$ and $\forall i,b_{i}\geq 0$. We prove that if $\mathsf{dist}(a,q)/\mathsf{dist}(b,q)\geq 1+\epsilon$,
@@
Case 2: $q^{(1)}\leq q^{\prime(1)}\leq 0$, the argument is similar.
D.2 A Data Structure for $n$ Points
In this section, we prove Theorem D.1. Let $C=\{c_{1},c_{2},\dots c_{n}\}\subseteq\mathbb{R}^{d}$ denote the set of input points and let $q\in\mathbb{R}^{d}$ denote the query point. Our goal is to find an approximate nearest neighbor of $q$ in the set $C$. The argument used here is from [1], but we still present it to be self-contained.
@@ D.2 A Data Structure for $n$ Points
For all $(c_{i},c_{j})\in\binom{C}{2}$, we run the preprocessing steps of the algorithm guaranteed by Lemma D.2, where we set $a\leftarrow c_{i}$ and $b\leftarrow c_{j}$.
@@
We first cast the nearest neighbor problem in the language of finding the minimum value in an array of length $n$ using only imperfect comparison queries as follows. Consider an array of length $n$ whose values are $\mathsf{dist}(q,c_{i})$ for $i\in[n]$. Although we do not have access to these values directly, Lemma D.2 guarantees an algorithm that, given two indices $i,j\in[n]$, with probability at least $1-\delta$, returns the symbol $>$ if $\mathsf{dist}(q,c_{i})>(1+\epsilon)\mathsf{dist}(q,c_{j})$ and returns the symbol $<$ if $(1+\epsilon)\mathsf{dist}(q,c_{i})<\mathsf{dist}(q,c_{j})$. In other words, we have access to a comparison operator for the entries of the array such that the comparator works correctly with high probability, if the entries are more than a factor of $1+\epsilon$ apart. We use the following recursive algorithm, called initially with $S$ set to $[n]$, can be used to find an approximate minimum in this case.
@@
    If $|S|=2$, run the algorithm guaranteed by Lemma D.2 and return the index of the smaller value. Note that this step corresponds to the use of the imperfect comparison operator.
@@
      For each pair $(m_{i},m_{j})\in\binom{M}{2}$, run the algorithm guaranteed by Lemma D.2.
@@
The recursion depth of the algorithm is $\log\log n$ by virtue of the recurrence $t(n)=t(\sqrt{n})+1$. The total number of calls to the query algorithm of Lemma D.2 at each recursion level is $\Theta(n)$ and hence, overall, this 2-point query algorithm is invoked $O(n\log\log n)$ times.
@@
We first condition on the good event that every call to the query algorithm guaranteed by Lemma D.2 is correct. This event happens with probability at least $1-\Theta(n\log\log n)\cdot\delta$. Next, we argue about the deterioration in the approximation guarantee. The main idea is that Step 2(d) and 2(e) together return a $(1+\epsilon)^{2}$-approximate nearest neighbor of $q$ in the set $M$, given our conditioning. Let $m^{*}\in M$ denote the nearest neighbor of $q$ in $M$. Assume for the sake of contradiction that the point $m\in M$ returned in Step 2(e) is such that $\mathsf{dist}(q,m)>(1+\epsilon)^{2}\cdot\mathsf{dist}(q,m^{*})$. By the guarantee of our imperfect comparison operator, $m$ does not "win" against $m^{*}$ in the 2-point comparison game. Consider now an $m^{\prime\prime}\in M$ that $m$ wins against in the 2-point comparison game. If $\mathsf{dist}(m^{\prime\prime},q)>(1+\epsilon)\cdot\mathsf{dist}(m,q)$, then it must also be the case that $m^{*}$ wins against $m^{\prime\prime}$ in the comparison game. Even if $\mathsf{dist}(m^{\prime\prime},q)\leq(1+\epsilon)\cdot\mathsf{dist}(m,q)$ and $m$ wins against $m^{\prime\prime}$ because of the do not care condition, the distance of $m^{\prime\prime}$ to $q$ is still a $(1+\epsilon)$ factor more than the distance of $m^{*}$ to $q$. Thus, $m^{*}$ one more 2-point comparison game as compared to $m$ and this is a contradiction. Since we incur a loss of a factor of $(1+\epsilon)^{2}$ in each recursion level, overall, the loss in approximation guarantee is $(1+\epsilon)^{2\log\log n}$.
Proof D.8 (Proof of Theorem D.1).
@@
To get our data structure, we run each invocation of the algorithm of Lemma D.2 above (i.e., in Min-Finding) with $\epsilon=\epsilon^{\prime}/\log\log n$ and $\delta=\delta^{\prime}/\Theta(n\log\log n)$, we can get a $(1+\epsilon^{\prime})$-approximate nearest neighbor data structure with success probability at least $1-\delta^{\prime}$.